OpenStudy (anonymous):

find the slope of the of the curve. y = 1 / (x-1) at x = 2

OpenStudy (anonymous):

Slope of a function is it's first derivative. So: m=dy/dx=d(x-1)^-1 /dx=-1(x-1)^-2 m=-1/(x-1)^2 Plug 2 into this eq to find m at x=2

OpenStudy (anonymous):

ok i need help finding the derivative

OpenStudy (anonymous):

i use the \[\lim_{h \rightarrow 0} \frac{ f(x+h) - f(x)}{ h} \] right?

OpenStudy (anonymous):

Correct, taht's the definition of derivative. Or simply: \[\frac{du^n}{dx}=n u^{n-1}*\frac{du}{dx}\]

OpenStudy (across):

In this case, you have to use something called the quotient rule:\[\frac d{dx}\left[\frac{g(x)}{h(x)}\right]=\frac{g'(x)h(x)-g(x)h'(x)}{h^2(x)}.\]